In how many ways can a research team of 5 members be formed from a group of 10 scientists consisting of 3 chemists and 7 physicists if the team must include exactly 2 chemists?
Answer by jim_thompson5910(35256)

There are 3*2 = 6 ways to pick 2 chemists assuming order matters. Order doesn't matter though which means you have 6/2! = 6/2 = 3 ways to pick 2 chemists.

Put another way, there are 3 ways to NOT pick a chemist.

-------------------------------------------------------

You have 5 slots. Two of these slots must be chemists. So you really have 5-2 = 3 slots left.

You cannot have another chemist (that last third chemist) on the team because you must have *exactly* 2 chemists. No more and no less.

So you have 3 slots left and 7 people to choose from to fill those three slots.

-------------------------------------------------------

There are 7*6*5 = 42*5 = 210 ways to pick 3 people from a pool of 7

This is if order mattered.

However, order doesn't matter, so you really have 210/3! = 210/6 = 35 ways to pick 3 people from a pool of 7.

-------------------------------------------------------

To sum up what we have so far:

We have 3 ways to pick two chemists
We have 35 ways to pick three physicists

So there are 3*35 = 105 ways to pick a team of five members.

Again, order does not matter.